OpenStudy (anonymous):

find lim as x approaches 1(x^3-1/x-1) I got does not exist because I did a sin chart and got -infinity and positive infinity. My book says the answer is 3. How would I do this problem? I plugged in one to the sin chart but got it wrong help please.

jimthompson5910 (jim_thompson5910):

hint: x^3 - 1 = x^3 - 1^3 use the formula a^3 - b^3 = (a-b)(a^2 + ab + b^2)

OpenStudy (anonymous):

When ever you have a polynomial over a polynomial like this, you have to factor. It should be your first instinct.